### 13 Mar 2026 v3

@@ Abstract.
We generalize the duality principle for the $\Gamma$-convergence of proper convex lower semicontinuous functions, which was previously known only in separable reflexive Banach spaces. Aditionally, We extend this duality to the broader class of weakly compactly generated (WCG) Banach spaces. Under a suitable hypotheses, we show that $\Gamma$-convergence in the strong topology is equivalent to $\Gamma$-convergence of the Fenchel conjugates in the weak topology. We further prove that this duality is equivalent to the graphical convergence of the associated subdifferentials with respect to the product topology given by the norm on the primal space and the weak topology on the dual. The extension to the WCG framework allows the inclusion of every separable space and every reflexive space, as well as important non-reflexive that might not be separable, such as $L^{1}(\mu)$ for an arbitrary $\sigma$-finite measure. As an application, we derive dual characterizations of the $\Gamma$-convergence of convex integral functionals on $L^{p}$ spaces. The proof of the main result relies on a separable reduction approach in WCG spaces using modern tools of Banach space theory, which, as a byproduct, provides new insight into the internal structure of WCG Banach spaces, in a spirit closely related to the structural phenomena underlying the theory of Asplund spaces [Cúth and Fabian (2016)].
@@ 1. Introduction
Despite these developments, the duality principle for $\Gamma$-convergence has remained confined, to the best of our knowledge, to reflexive separable Banach spaces. The role of reflexivity and separability in this context is not merely technical: both enter naturally in the classical proofs through compactness and metrizability arguments. It is therefore natural to ask whether these assumptions are intrinsic to the duality mechanism itself.
The main result of this paper shows that they are not, and that the natural dual framework for $\Gamma$-convergence of convex functions is given by the strong topology in the primal space and the weak topology in the dual. Within this framework, we establish a Fenchel duality principle for $\Gamma$-convergence in weakly compactly generated Banach spaces.
Our results also provide an answer to a question raised by Fitzpatrick and Lewis in [48], concerning the relation between the convergence of convex functions and the convergence of their Fenchel conjugates, moreover, the result covered by [48] is a particular case of our main result. Fitzpatrick question raised at a time where continuity of the Fenchel conjugate and the polar map, and generalizations of Mosco convergence to non-reflexive spaces, as well as topologies for closed (or closed and convex) sets and convex functions, were extensively studied (see, e.g., [15, 21, 22, 4, 43, 5, 36, 65, 20, 72, 19, 18, 26, 28, 17, 24, 66]).
@@ 6. Some Preliminary Results
In this section, we provide a proof of Theorem 4.1. The result is first established for separable Banach spaces and then extended to general WCG Banach spaces by means of separable reduction techniques, combined with the use of projectional skeletons.

### 16 Mar 2026 v4

@@ Abstract.
We generalize the duality principle for the $\Gamma$-convergence of proper convex lower semicontinuous functions, which was previously established only in separable reflexive Banach spaces. Moreover, we extend this principle to the broader class of weakly compactly generated (WCG) Banach spaces. Under a suitable hypothesis, we show that $\Gamma$-convergence in the norm topology is equivalent to $\Gamma$-convergence of the Fenchel conjugates in the weak topology. We further prove that this duality is equivalent to the graphical convergence of the associated subdifferentials with respect to the product topology given by the norm on the primal space and the weak topology on the dual. The WCG setting encompasses all separable and all reflexive Banach spaces, as well as important non-reflexive spaces which may fail to be separable, such as $L^{1}(\mu)$ for an arbitrary $\sigma$-finite measure. In the reflexive case, the generalization recovers the classical results without assuming separability. As an application, we derive dual characterizations of the $\Gamma$-convergence of convex integral functionals on $L^{p}$ spaces. The proof relies on a separable reduction argument in WCG spaces based on modern tools from Banach space theory. As a byproduct, our approach provides additional insight into the structure of WCG spaces, in a direction reminiscent of phenomena arising in the theory of Asplund spaces [Cúth and Fabian (2016)].
@@ 1. Introduction
Despite these developments, the duality theory for $\Gamma$-convergence has classically been confined, to the best of our knowledge, to separable reflexive Banach spaces,a restriction which arises from genuine intrinsic technical obstacles in the available approaches. It is therefore natural to ask whether these assumptions are inherent in the duality mechanism itself.
In this work, relying on more recent advances in Banach space theory, we show that these constraints can be removed: separability and reflexivity are not required. Even in the reflexive setting separability is not necessary, and the duality framework can in fact be extended to the broader class of weakly compactly generated spaces.
Furthermore, the main result of this paper shows that the natural dual framework for $\Gamma$-convergence of convex functions is given by the strong topology in the primal space and the weak topology in the dual. Within this framework, we establish a Fenchel duality principle for $\Gamma$-convergence in weakly compactly generated Banach spaces.
Our results also provide an answer to a question raised by Fitzpatrick and Lewis in [48], concerning the relation between the convergence of convex functions and the convergence of their Fenchel conjugates, moreover, the result covered by [48] is a particular case of our main result. Fitzpatrick’s question raised at a time where continuity of the Fenchel conjugate and the polar map, and generalizations of Mosco convergence to non-reflexive spaces, as well as topologies for closed (or closed and convex) sets and convex functions, were extensively studied (see, e.g., [15, 21, 22, 4, 43, 5, 36, 65, 20, 72, 19, 18, 26, 28, 17, 24, 66]).
@@ 6. Some Preliminary Results
In this section, we provide a proof of Theorem 4.1. The result is first established for separable Banach spaces and then extended to general WCG Banach spaces by means of separable reduction techniques, supported by the use of projectional skeletons.